# Offline changepoint localization using a matrix of conformal p-values

### v4

@@ 1 Introduction
This work focuses on localizing a single changepoint; one possible extension (which we do not explore here) is to use a variant of our algorithm to localize multiple changepoints. We begin by reviewing prior work on changepoint localization and conformal approaches to change detection in Section 2. We introduce the MCP algorithm in Section 3. We then show how to construct confidence sets for the changepoint in Section 4 and prove the validity of the resulting confidence sets. Next, we discuss how the score function should be chosen in Section 5. Finally, we apply the MCP algorithm on a variety of synthetic and real-world datasets in Section 6 to demonstrate its effectiveness and flexibility.
@@ 2 Related work
Here, we give a review of several classical methods from changepoint analysis; some further details can be found in [25]. Traditional changepoint analysis has largely been dominated by parametric approaches based on the generalized likelihood ratio; however, these parametric methods often require knowing the pre-change and post-change distributions to achieve optimality, and often only give point estimators without associated confidence sets.
@@
Methods such as CUSUM [15] and conformal martingales [29] are used for quickest (sequential) changepoint detection. One cannot use them to get confidence sets for the location of the changepoint (sometimes called changepoint localization). In fact, there is only one very recent work of [19] that attempts to provide confidence sets for changepoints after stopping a sequential detection procedure, but they make parametric assumptions, as well as assume that the pre- and post-change distributions are in known non-overlapping classes. We do not need such assumptions, and methods in their preprint are not applicable to our assumption-lean setup.
Nonparametric changepoint localization
@@
There is previous work in nonparametric changepoint localization specifically for a change in the mean, but they rely on unspecified hyperparameters or are only asymptotically valid. Furthermore, many interesting changepoints do not involve a mean change (e.g., they may involve a change in the variance, modality, etc.). In [26], the authors provide a CI for the changepoint, but the confidence interval in their Proposition 4 has several unknown constants (only their existence is ensured), and they do not provide a method to explicitly compute them (which is perhaps why the paper has no experiments). As a result, we could not implement their CI in practice, even in their special case of mean change. The confidence set in [5] works only for mean changes in univariate data, is only asymptotically valid, and relies on a computationally expensive bootstrap procedure.
@@
The SMUCE estimator introduced in [10] provides an asymptotically valid confidence set in exponential family regression models. The paper [31] produces a confidence set for high-dimensional regression problems under many technical assumptions, and the result in [9] can be used to construct asymptotic CIs for a Gaussian mean change. However, note that all of these methods lack finite-sample validity and operate under far more stringent assumptions than our MCP algorithm.
@@
Rank-based nonparametric tests, such as those developed by [16] and [18], offer distribution-free tests about whether there was a changepoint in a given dataset, but these methods do not produce a confidence set for the changepoint on rejection and suffer from lack of statistical power without additional assumptions. In contrast, MCP addresses these limitations by allowing the score function to be learned in a way that achieves nontrivial power, while also providing finite-sample valid confidence sets for changepoint location under minimal distributional assumptions.
@@
Existing methods for multiple changepoint analysis (which typically only yield point estimators) often rely on the so-called ”Isolate-Detect” paradigm, using a sliding window or segmentation to reduce to the single changepoint case (for instance, see Section 5.2 of [25] or [1]. In this sense, our algorithm is an important contribution to the changepoint analysis literature, since (after isolation of the changepoint) it provides a method for producing a confidence set; future work may include using our algorithm as a subroutine in one of these ”isolate-detect” procedures to localize multiple changepoints.
Conformal prediction for change detection.
The application of conformal prediction to changepoint problems is relatively recent compared to traditional approaches. Conformal prediction, as surveyed in [22], provides a framework for constructing valid prediction regions with distribution-free guarantees under minimal statistical assumptions. While originally developed for supervised learning tasks, conformal methods have been extended to changepoint detection.
Early work on connecting conformal prediction to changepoint detection focused primarily on testing for exchangeability violations rather than precise localization. [29] proposed a general approach for testing the exchangeability assumption in an online setting using conformal martingales, laying important groundwork for changepoint detection. The conformal martingale approach is a special case of the e-detector framework laid out in [23], which provides an extremely general and powerful method for sequential change detection. While most prior work using conformal prediction for change detection is in the sequential (online) setting, here we study the localization problem in the offline setting; we would like a confidence interval for the changepoint.
Building on this foundation, [28] and [30] introduced conformal test martingales specifically designed for changepoint detection; for them, the motivation was to determine when the data generating distribution changes and a predictive algorithm needs to be retrained. They introduced conformal versions of the CUSUM and Shiryaev-Roberts procedures, which control false alarm rates through martingale properties. However, the primary emphasis remained on detection rather than localization of the changepoint.
[27] developed a more computationally efficient conformal test martingale, called the inductive conformal martingale, for change detection. Furthermore, they provided conformity measures and betting functions tailored specifically for change detection. However, their method does not provide formal guarantees for the localization task. More recently, [14] investigated conformal changepoint detection under the assumption that the data generating distribution is continuous, showing that the conformal martingale is statistically efficient but without addressing the question of confidence interval construction for the changepoint location.
Despite these recent advances, existing conformal changepoint methods focus on detection (whether a change occurred) rather than localization (precise estimation of where it occurred). The primary novelty of our approach lies in leveraging conformal prediction techniques to construct valid confidence sets for the changepoint location, as well as allowing for the score function to be learned from the data. By bridging the gap between conformal prediction and classical changepoint localization using two-sample testing, our approach opens new possibilities for reliable changepoint analysis in complex, high-dimensional data settings where classical methods fail.
@@ 3 The matrix of conformal p-values (MCP) algorithm
In this section, we discuss our MCP algorithm for conformal changepoint localization in technical detail. We begin by defining score functions; for more details on conformal prediction, see [22]. In the following definition, we use $\llbracket\mathcal{X}^{m}\rrbracket$ to denote the set of unordered bags of $m$ data points (which may contain repetitions); such an unordered bag is often called a multiset. We will denote an element of $\llbracket\mathcal{X}^{m}\rrbracket$ by $\llbracket Y_{1},\dots,Y_{m}\rrbracket$, where the $Y_{i}$ are $\mathcal{X}$-valued random variables.
@@
A family of score functions is a list $((s_{rt})_{r=1}^{n-1})_{t=1}^{n}$ of functions $s_{rt}:\mathcal{X}\times\llbracket\mathcal{X}^{r}\rrbracket\times\mathcal{X}^{n-t}\to\mathbb{R}$. Each element in a family of score functions is called a score function.
Note that a score function is intuitively intended to be a pre-processing transformation intended to separate the pre-change and post-change data points by projecting them into one dimension. In particular, the score function can be learned in any way that uses its second argument exchangeably, while its third argument can be used non-exchangeably. The goal of our algorithms is to simultaneously test the null hypotheses that a change occurs at time $t\in[n-1]$:
@@
$\mathcal{C}_{1-\alpha}=\left\{t\in[n-1]:p_{t}>\alpha\right\},$
Since $\mathcal{H}_{0t}$ is only a test of exchangeability on either side of the changepoint, the above confidence interval is not valid if there is no changepoint in the data. However, we describe a method to pre-test the data for exchangeability in Appendix D, which we empirically observe does not hurt the performance of our algorithms. On the other hand, if a point estimator for the changepoint is desired, we can output the estimate $\hat{\xi}=\arg\max_{t\in[n-1]}\;p_{t}$.
As described in Section 1, we assume that $\mathbb{P}_{0}$ and $\mathbb{P}_{1}$ are exchangeable. Then, we have the following algorithm for changepoint localization, which we call the MCP (matrix of conformal p-values) algorithm. In the following algorithm, let $u$ denote the cumulative distribution of a $\mathrm{Unif}(0,1)$ distribution. Furthermore, recall that the Kolmogorov-Smirnov (KS) distance between two cumulative distributions $F$ and $G$ is given by $\mathrm{KS}(F,G)=\sup_{z\in\mathbb{R}}\;\lvert F(z)-G(z)\rvert$. Our algorithm is motivated by the classical Kolmogorov-Smirnov test for goodness of fit. Note that our algorithm does not specially depend on the choice of the KS distance; variants are in Appendix C.
Input: $(X_{t})_{t=1}^{n}$ (dataset), $((s_{rt}^{(0)})_{r=1}^{n-1})_{t=1}^{n}$ and $((s_{rt}^{(1)})_{r=1}^{n-1})_{t=1}^{n}$ (left and right score function families)
@@
1
2for $t\in[n-1]$ do
3
4 for $r$ in $(1,\dots,t)$ do
5    for $j$ in $(1,\dots,r)$ do
6       $\kappa_{rj}^{(t)}\leftarrow s_{rt}^{(0)}(X_{j}\,;\,\llbracket X_{1},\dots,X_{r}\rrbracket,(X_{t+1},\dots,X_{n}))$
7      end for
8
9    $p_{r}^{(t)}\leftarrow\frac{1}{r}\sum_{j=1}^{r}\left(\mathbf{1}_{\kappa_{rj}^{(t)}>\kappa_{rr}^{(t)}}+\theta_{r}^{(t)}\,\mathbf{1}_{\kappa_{rj}^{(t)}=\kappa_{rr}^{(t)}}\right)$, where $\theta_{r}^{(t)}\sim\mathrm{Unif}(0,1)$
10   end for
11
12 for $r$ in $(n,\dots,t+1)$ do
13    for $j$ in $(r+1,\dots,n)$ do
14       $\kappa_{rj}^{(t)}\leftarrow s_{n-r,\,n-t}^{(1)}(X_{j}\,;\,\llbracket X_{r+1},\dots,X_{n}\rrbracket,(X_{1},\dots,X_{t}))$
15      end for
16    $p_{r}^{(t)}\leftarrow\frac{1}{n-r+1}\sum_{j=r}^{n}\left(\mathbf{1}_{\kappa_{rj}^{(t)}>\kappa_{rr}^{(t)}}+\theta_{r}^{(t)}\,\mathbf{1}_{\kappa_{rj}^{(t)}=\kappa_{rr}^{(t)}}\right)$, where $\theta_{r}^{(t)}\sim\mathrm{Unif}(0,1)$
17   end for
18
19 $\hat{F}_{0}(z):=\frac{1}{t}\sum_{r=1}^{t}\mathbf{1}_{p_{r}^{(t)}\leq z}$, $\hat{F}_{1}(z):=\frac{1}{n-t}\sum_{r=t+1}^{n}\mathbf{1}_{p_{r}^{(t)}\leq z}$
20 $W_{t}^{(0)}\leftarrow\sqrt{t}\;\mathrm{KS}(\hat{F}_{0},u)$, $W_{t}^{(1)}\leftarrow\sqrt{n-t}\;\mathrm{KS}(\hat{F}_{1},u)$
21  Use Algorithm 2 to map $(W_{t}^{(0)},W_{t}^{(1)})$ to left and right p-values $(p_{t}^{\mathrm{left}},p_{t}^{\mathrm{right}})$
22  Combine $p_{t}^{\mathrm{left}}$ and $p_{t}^{\mathrm{right}}$ into a single p-value $p_{t}$ (Section 4.2)
23 end for
24
[…]
@@
Here is some intuition for the MCP algorithm (Algorithm 1). First, the algorithm splits the data into left and right halves at $t\in[n-1]$; we focus on the left half, since the right half is analogous. The goal of MCP is first to create p-values $(p_{r}^{(t)})_{r=1}^{t}$ which are independent. For each $r\in[t]$, we map the data points $(X_{j})_{j=1}^{r}$ through a score function into exchangeable one-dimensional statistics $(\kappa_{rj}^{(t)})_{j=1}^{r}$ (algorithm 1). Then, the (normalized) rank statistics of the scores are sequentially computed on the left and right halves (algorithm 1). If there was no change in the left segment, the ranks for that segment would each be distributed like $\mathrm{Unif}(0,1)$. In particular, as $t\to\infty$, we know by the Glivenko-Cantelli theorem that $\sup_{z\in\mathbb{R}}\;\lvert\hat{F}_{0}(z)-u(z)\rvert\to 0$.
Note that because $W_{t}^{(0)}$ only depends on the normalized ranks $p_{r}^{(t)}$ for $r\in[t]$, it is a distribution-free statistic under $\mathcal{H}_{0t}$ and we can use it for hypothesis testing. Essentially, $W_{t}^{(0)}$ measures how far the normalized ranks in the left and right halves of the ordered list are from being uniform. We can convert this statistic to a p-value by a hypothesis test (algorithm 1) which can be combined with the p-value obtained from the right half for each $t$ (algorithm 1). Finally, the test can be inverted to form a confidence set by the Neyman construction (algorithm 1).
To motivate the normalization by $\sqrt{t}$ in the definition of $W_{t}^{(0)}$, recall that we have samples $p_{r}^{(t)}\sim\operatorname{Unif}(0,1)$ for $r\in[t]$ and let $(\varphi_{z})_{z\geq 0}$ denote the Brownian bridge ($\varphi_{z}=B_{z}-zB_{1}$, where $(B_{z})_{z\geq 0}$ is a standard Brownian motion on $\mathbb{R}$). Then, a central limit theorem by [7] implies that $\sqrt{t}\,\left(\sup_{z\in\mathbb{R}}\;\lvert\hat{F}_{0}(z)-u(z)\rvert\right)=\sqrt{t}\;\mathrm{KS}(\hat{F}_{0},u)\overset{d}{\longrightarrow}\sup_{z\in[0,1]}\;\lvert\varphi_{z}\rvert$. Notice that normalizing by $\sqrt{t}$ stabilizes the KS distance, and we have an explicit form for the asymptotic distribution.
Algorithm 1 effectively computes the p-values $p_{r}^{(t)}$, which can be written as an $n\times(n-1)$ matrix; we call this the matrix of conformal p-values (MCP):
$\mathrm{MCP}:=\begin{bmatrix}p_{1}^{(1)}&p_{1}^{(2)}&\cdots&p_{1}^{(n-1)}\\ p_{2}^{(1)}&p_{2}^{(2)}&\cdots&p_{2}^{(n-1)}\\ \vdots&\vdots&\vdots&\vdots\\ p_{n}^{(1)}&p_{n}^{(2)}&\cdots&p_{n}^{(n-1)}\end{bmatrix}$.
For each candidate changepoint $t\in[n-1]$ we construct $n$ p-values for $\mathcal{H}_{0t}$ (corresponding to a column in the MCP), where each p-value corresponds to one of the data points. Recall from Algorithm 1 that these p-values are constructed by considering the randomized rank of that point’s score, relative to the other points on the same side. Focusing on the $t$th row of the MCP, we then refine these p-values as in Figure 2.
@@
Figure 2: Refinement of the p-values in the $t$th row of the matrix of conformal p-values (MCP) into the discrepancy statistics $W_{t}^{(0)}$ and $W_{t}^{(1)}$.
Focus on a particular index $t\in[n-1]$. Under $\mathcal{H}_{0t}$, the scores $(\kappa_{rj}^{(t)})_{j=1}^{r}$ are exchangeable for $r\leq t$, meaning that the p-values $(p_{r}^{(t)})_{r=1}^{t}$ are independent. In Section 4, we discuss the subroutines of the MCP algorithm in greater depth. We have the following finite-sample coverage guarantee.
Theorem 3.1 (Coverage guarantee (empirical test)).
$\mathbb{P}\left(\xi\in\mathcal{C}_{1-\alpha}\right)\geq 1-\alpha.$
4 Constructing confidence sets for the changepoint
In this section, we describe several methods for constructing confidence sets for the changepoint based on the Neyman construction, which inverts the hypothesis test of $\mathcal{H}_{0t}:\xi=t$ against $\mathcal{H}_{1t}:\xi\neq t$. Next, we will produce two p-values for this test, and then combine them into one.

### v5

@@ 1 Introduction
This work focuses on localizing a single changepoint; one possible extension (which we do not explore here) is to use a variant of our algorithm to localize multiple changepoints. We begin by reviewing prior work on changepoint localization and conformal approaches to change detection in Section 2. We introduce the MCP algorithm in Section 3. We then detail several of the subroutines of the MCP algorithm in Section 4 and prove the validity of the resulting confidence sets. Next, we discuss how the score function should be chosen in Section 5. Finally, we apply the MCP algorithm on a variety of synthetic and real-world datasets in Section 7 to demonstrate its effectiveness and flexibility.
@@ 2 Related work
Here, we give a review of several classical methods from changepoint analysis; some further details can be found in [truong2020selective]. Traditional changepoint analysis has largely been dominated by parametric approaches based on the generalized likelihood ratio; however, these parametric methods often require knowing the pre-change and post-change distributions to achieve optimality, and often only give point estimators without associated confidence sets.
@@
Methods such as CUSUM [page1955test] and conformal martingales [vovk2003testing] are used for quickest (sequential) changepoint detection. One cannot use them to get confidence sets for the location of the changepoint (sometimes called changepoint localization). In fact, there is only one very recent work of [saha2025post] that attempts to provide confidence sets for changepoints after stopping a sequential detection procedure, but they make parametric assumptions, as well as assume that the pre- and post-change distributions are in known non-overlapping classes. We do not need such assumptions, and methods in their preprint are not applicable to our assumption-lean setup.
Conformal prediction for change detection.
The application of conformal prediction to changepoint problems is relatively recent compared to traditional approaches. Conformal prediction, as surveyed in [shafer2008tutorial], provides a framework for constructing valid prediction regions with distribution-free guarantees under minimal statistical assumptions. While originally developed for supervised learning tasks, conformal methods have been extended to changepoint detection.
In contrast to traditional approaches to quickest changepoint detection (detecting whether a change occurred) using conformal martingales [vovk2003testing], our method allows for localization (precise estimation of where the changepoint occurred). Our main technical contribution is to leverage conformal p-values constructed online from the data in the conformal martingale framework; these p-values are known to be i.i.d. and distributed as $\mathrm{Unif}(0,1)$ if no change has occurred yet. By a two-sample test, we are then able to localize the changepoint. Furthermore, in contrast to other conformal approaches to changepoint analysis, we allow for the score function to be learned from the data and prove a novel conformal Neyman-Pearson lemma guiding the choice of these score functions.
Early work on connecting conformal prediction to changepoint detection focused primarily on testing for exchangeability violations rather than precise localization. [vovk2003testing] proposed a general approach for testing the exchangeability assumption in an online setting using conformal martingales, laying important groundwork for changepoint detection. The conformal martingale approach is a special case of the e-detector framework laid out in [shin2022detectors], which provides an extremely general and powerful method for sequential change detection. While most prior work using conformal prediction for change detection is in the sequential (online) setting, here we study the localization problem in the offline setting; we would like a confidence interval for the changepoint.
Building on this foundation, [vovk2021testing] and [vovk2021retrain] introduced conformal test martingales specifically designed for changepoint detection; for them, the motivation was to determine when the data generating distribution changes and a predictive algorithm needs to be retrained. They introduced conformal versions of the CUSUM and Shiryaev-Roberts procedures, which control false alarm rates through martingale properties. However, the primary emphasis remained on detection rather than localization of the changepoint.
[volkhonskiy2017inductive] developed a more computationally efficient conformal test martingale, called the inductive conformal martingale, for change detection. Furthermore, they provided conformity measures and betting functions tailored specifically for change detection. However, their method does not provide formal guarantees for the localization task. More recently, [nouretdinov2021conformal] investigated conformal changepoint detection under the assumption that the data generating distribution is continuous, showing that the conformal martingale is statistically efficient but without addressing the question of confidence interval construction for the changepoint location.
By bridging the gap between conformal prediction and classical changepoint localization using two-sample testing, our approach opens new possibilities for reliable changepoint analysis in complex, high-dimensional data settings where classical methods fail.
Changepoint localization
@@
There is previous work in changepoint localization specifically for a change in the mean, but they rely on unspecified hyperparameters or are only asymptotically valid. Furthermore, many interesting changepoints do not involve a mean change (e.g., they may involve a change in the variance, modality, etc.). In [verzelen2023optimal], the authors provide a CI for the changepoint, but the confidence interval in their Proposition 4 has several unknown constants (only their existence is ensured), and they do not provide a method to explicitly compute them (which is perhaps why the paper has no experiments). As a result, we could not implement their CI in practice, even in their special case of mean change. The confidence set in [cho2022bootstrap] works only for mean changes in univariate data, is only asymptotically valid, and relies on a computationally expensive bootstrap procedure.
@@
The SMUCE estimator introduced in [frick2014multiscale] provides an asymptotically valid confidence set in exponential family regression models. The paper [xu2024change] produces a confidence set for high-dimensional regression problems under many technical assumptions, and the result in [fotopoulos2010exact] can be used to construct asymptotic CIs for a Gaussian mean change. However, note that all of these methods lack finite-sample validity and operate under far more stringent assumptions than our MCP algorithm.
[carlstein1988nonparametric, lee1996nonparametric, zou2014nonparametric] study general nonparametric changepoint problems, but only provide consistent point estimators of the changepoint. [dumbgen1991asymptotic] provides a method to produce confidence sets in the general nonparametric setting, but this method is only asymptotically valid, requires a computationally expensive bootstrap procedure, and relies on several complex theoretical assumptions which are difficult to verify in practice. When instantiated to an exponential family model, the method of [dumbgen1991asymptotic] provides an explicit formula for the confidence set matching that of [siegmund1988confidence].
@@
Rank-based nonparametric tests, such as those developed by [pettitt1979non] and [ross2012two], offer distribution-free tests about whether there was a changepoint in a given dataset, but these methods do not produce a confidence set for the changepoint on rejection and suffer from lack of statistical power without additional assumptions. In contrast, MCP addresses these limitations by allowing the score function to be learned in a way that achieves nontrivial power, while also providing finite-sample valid confidence sets for changepoint location under minimal distributional assumptions.
@@
Existing methods for multiple changepoint analysis (which typically only yield point estimators) often rely on the so-called ”Isolate-Detect” paradigm, using a sliding window or segmentation to reduce to the single changepoint case (for instance, see Section 5.2 of [truong2020selective] or [anastasiou2022detecting]). In this sense, our algorithm is an important contribution to the changepoint analysis literature, since (after isolation of the changepoint) it provides a method for producing a confidence set.
@@ 3 The matrix of conformal p-values (MCP) algorithm
In this section, we discuss our MCP algorithm for conformal changepoint localization in technical detail. We begin by defining score functions; for more details on conformal prediction, see [shafer2008tutorial]. In the following definition, we use $\llbracket\mathcal{X}^{m}\rrbracket$ to denote the set of unordered bags of $m$ data points (which may contain repetitions); such an unordered bag is often called a multiset. We will denote an element of $\llbracket\mathcal{X}^{m}\rrbracket$ by $\llbracket Y_{1},\dots,Y_{m}\rrbracket$, where the $Y_{i}$ are $\mathcal{X}$-valued random variables.
@@
A family of score functions is a list $((s_{rt})_{r=1}^{n})_{t=1}^{n}$ of functions $s_{rt}:\mathcal{X}\times\llbracket\mathcal{X}^{r}\rrbracket\times\mathcal{X}^{n-t}\to\mathbb{R}$. Each element in a family of score functions is called a score function.
Note that a score function is intuitively intended to be a pre-processing transformation intended to separate the pre-change and post-change data points by projecting them into one dimension. In particular, the score function can be learned in any way that uses its second argument exchangeably, while its third argument can be used non-exchangeably. For intuition, we provide several examples of valid score functions in Appendix I and at the end of Section 5. The goal of our algorithms is to simultaneously test the null hypotheses that a change occurs at time $t\in[n]$:
@@
$\mathcal{C}_{1-\alpha}=\left\{t\in[n]:p_{t}>\alpha\right\},$
If $n$ is present in the resulting confidence set, the practitioner should interpret this to mean that there was possibly no changepoint present in the dataset. On the other hand, if a point estimator for the changepoint is desired, we can output the estimate $\hat{\xi}=\arg\max_{t\in[n]}\;p_{t}$.
As described in Section 1, we assume that $\mathbb{P}_{0}$ and $\mathbb{P}_{1}$ are exchangeable. Then, we have the following algorithm for changepoint localization, which we call the MCP (matrix of conformal p-values) algorithm. In the following algorithm, let $u$ denote the cumulative distribution of a $\mathrm{Unif}(0,1)$ distribution. Furthermore, recall that the Kolmogorov-Smirnov (KS) distance between two cumulative distributions $F$ and $G$ is given by $\mathrm{KS}(F,G)=\sup_{z\in\mathbb{R}}\;\lvert F(z)-G(z)\rvert$. Our algorithm is motivated by the classical Kolmogorov-Smirnov test for goodness of fit. Note that our algorithm does not specially depend on the choice of the KS distance; variants are in Appendix D.
Input: $(X_{t})_{t=1}^{n}$ (dataset), $((s_{rt}^{(0)})_{r=1}^{n})_{t=1}^{n}$ and $((s_{rt}^{(1)})_{r=1}^{n})_{t=1}^{n}$ (left and right score function families)
@@
for $t\in[n]$ do
 for $r$ in $(1,\dots,t)$ do
    for $j$ in $(1,\dots,r)$ do
       $\kappa_{rj}^{(t)}\leftarrow s_{rt}^{(0)}(X_{j}\,;\,\llbracket X_{1},\dots,X_{r}\rrbracket,(X_{t+1},\dots,X_{n}))$
      end for
    // compute per-anomaly p-values
    $p_{r}^{(t)}\leftarrow\frac{1}{r}\sum_{j=1}^{r}\left(\mathbf{1}_{\kappa_{rj}^{(t)}>\kappa_{rr}^{(t)}}+\theta_{r}^{(t)}\,\mathbf{1}_{\kappa_{rj}^{(t)}=\kappa_{rr}^{(t)}}\right)$, where $\theta_{r}^{(t)}\sim\mathrm{Unif}(0,1)$
   end for
 for $r$ in $(n,\dots,t+1)$ do
    for $j$ in $(r,\dots,n)$ do
       $\kappa_{rj}^{(t)}\leftarrow s_{n-r,\,n-t}^{(1)}(X_{j}\,;\,\llbracket X_{r+1},\dots,X_{n}\rrbracket,(X_{1},\dots,X_{t}))$
      end for
    // compute per-anomaly p-values
    $p_{r}^{(t)}\leftarrow\frac{1}{n-r+1}\sum_{j=r}^{n}\left(\mathbf{1}_{\kappa_{rj}^{(t)}>\kappa_{rr}^{(t)}}+\theta_{r}^{(t)}\,\mathbf{1}_{\kappa_{rj}^{(t)}=\kappa_{rr}^{(t)}}\right)$, where $\theta_{r}^{(t)}\sim\mathrm{Unif}(0,1)$
   end for
 $\hat{F}_{0}(z)\coloneqq\frac{1}{t}\sum_{r=1}^{t}\mathbf{1}_{p_{r}^{(t)}\leq z}$, $\hat{F}_{1}(z)\coloneqq\frac{1}{n-t}\sum_{r=t+1}^{n}\mathbf{1}_{p_{r}^{(t)}\leq z}$
 $W_{t}^{(0)}\leftarrow\sqrt{t}\;\mathrm{KS}(\hat{F}_{0},u)$, $W_{t}^{(1)}\leftarrow\sqrt{n-t}\;\mathrm{KS}(\hat{F}_{1},u)$
  Use Algorithm 2 to map $(W_{t}^{(0)},W_{t}^{(1)})$ to left and right p-values $(p_{t}^{\mathrm{left}},p_{t}^{\mathrm{right}})$
  Combine $p_{t}^{\mathrm{left}}$ and $p_{t}^{\mathrm{right}}$ into a per-candidate p-value $p_{t}$ (Section 4.2)
end for
$\mathcal{C}_{1-\alpha}\leftarrow\{t\in[n]:p_{t}>\alpha\}$
@@
Before providing further intuition for MCP, we first note that the case $t=n$ may proceed slightly differently in order to improve power.
Remark.
For the computation of $p_{n}$, we also recommend to compute the sequential ranks in reverse by
and calculate a “backward p-value” using a one-sample KS test from uniform using $(\overline{p}_{1},\dots,\overline{p}_{n})$. Then, the forward and backward p-values obtained from the respective KS tests can be combined using the Bonferroni method (Section 4.2). If the practitioner knows ahead of time that a changepoint exists in the dataset, the test for $\mathcal{H}_{0t}$ may be skipped in Algorithm 1 and the validity guarantee continues to hold.
Algorithm 1 effectively computes all of the per-anomaly p-values $p_{r}^{(t)}$, which can be written as an $n\times n$ matrix; we call this the matrix of conformal p-values (MCP):
$\mathrm{MCP{}}\coloneqq\begin{bmatrix}p_{1}^{(1)}&p_{1}^{(2)}&\cdots&p_{1}^{(n)}\\ p_{2}^{(1)}&p_{2}^{(2)}&\cdots&p_{2}^{(n)}\\ \vdots&\vdots&\vdots&\vdots\\ p_{n}^{(1)}&p_{n}^{(2)}&\cdots&p_{n}^{(n)}\end{bmatrix}$.
For each candidate changepoint $t\in[n]$ we construct $n$ per-anomaly p-values for $\mathcal{H}_{0t}$ (corresponding to a column in the MCP), where each per-anomaly p-value corresponds to one of the data points. Recall from Algorithm 1 that these p-values are constructed by considering the randomized rank of that point’s score, relative to the other points on the same side. Focusing on the $t$th column of the MCP, we then refine these p-values into left and right p-values as in Figure 2. Lastly, $p_{t}^{\mathrm{left}}$ and $p_{t}^{\mathrm{right}}$ are combined into the per-candidate p-value $p_{t}$, which we use to perform a hypothesis test of $\mathcal{H}_{0t}:\xi=t$ for a candidate changepoint. The computational complexity of the MCP algorithm is dependent on the choice of score function, and the base algorithm runs in $O(n^{2})$ time. In fact, most of the score functions that we suggest in Section 5 and Appendix I can be computed with an $O(n^{2})$ pre-processing step, so our algorithm typically has an overall quadratic complexity.
@@
Figure 2: Refinement of the per-anomaly p-values in the $t$th row of the matrix of conformal p-values (MCP) into left and right p-values.
Here is some further intuition for the MCP algorithm (Algorithm 1). First, the algorithm splits the data into left and right halves at $t\in[n]$; we focus on the left half, since the right half is analogous. The goal of MCP is first to create p-values $(p_{r}^{(t)})_{r=1}^{t}$ which are independent. For each $r\in[t]$, we map the data points $(X_{j})_{j=1}^{r}$ through a score function into exchangeable one-dimensional statistics $(\kappa_{rj}^{(t)})_{j=1}^{r}$ (algorithm 1). Then, the (normalized) rank statistics of the scores are sequentially computed on the left and right halves (algorithm 1). If there was no change in the left segment, the ranks for that segment would each be distributed like $\mathrm{Unif}(0,1)$. In particular, as $t\to\infty$, we know by the Glivenko-Cantelli theorem that $\sup_{z\in\mathbb{R}}\;\lvert\hat{F}_{0}(z)-u(z)\rvert\to 0$.
Note that because $W_{t}^{(0)}$ only depends on the per-anomaly p-value $p_{r}^{(t)}$ for $r\in[t]$, it is a distribution-free statistic under $\mathcal{H}_{0t}$ and we can use it for hypothesis testing. Essentially, $W_{t}^{(0)}$ measures how far the normalized ranks in the left and right halves of the ordered list are from being uniform. We can convert this statistic to a p-value by a hypothesis test (algorithm 1) which can be combined with the p-value obtained from the right half for each $t$ (algorithm 1). Finally, we compute $p_{n}$ by testing for exchangeability as in the remark following Algorithm 1 and the tests can simultaneously be inverted to form a confidence set by the Neyman construction (algorithm 1).
To motivate the normalization by $\sqrt{t}$ in the definition of $W_{t}^{(0)}$, recall that we have samples $p_{r}^{(t)}\sim\operatorname{Unif}(0,1)$ for $r\in[t]$ and let $(\varphi_{z})_{z\geq 0}$ denote the Brownian bridge ($\varphi_{z}=B_{z}-zB_{1}$, where $(B_{z})_{z\geq 0}$ is a standard Brownian motion on $\mathbb{R}$). Then, a central limit theorem by [donsker1951invariance] implies that $\sqrt{t}\,\left(\sup_{z\in\mathbb{R}}\;\lvert\hat{F}_{0}(z)-u(z)\rvert\right)=\sqrt{t}\;\mathrm{KS}(\hat{F}_{0},u)\overset{d}{\longrightarrow}\sup_{z\in[0,1]}\;\lvert\varphi_{z}\rvert$. Notice that normalizing by $\sqrt{t}$ stabilizes the KS distance, and we have an explicit form for the asymptotic distribution.
Under $\mathcal{H}_{0t}$, the scores $(\kappa_{rj}^{(t)})_{j=1}^{r}$ are exchangeable for $r\leq t$, meaning that the per-anomaly p-values are independent and uniform both to the left and right of $t$. We then aggregate the per-anomaly p-values to create the per-candidate p-value $p_{t}$, which allows us to test $\mathcal{H}_{0t}$. In Section 4, we discuss the subroutines of the MCP algorithm in greater depth. Note that Algorithm 1 is also symmetric under reflections of the data; if we reversed the input data as $\tilde{X}\coloneqq(X_{n},\dots,X_{1})$, the MCP algorithm would output $\tilde{\mathcal{C}}_{1-\alpha}\coloneqq n-\mathcal{C}_{1-\alpha}$, where $\mathcal{C}_{1-\alpha}$ is the output of Algorithm 1 on $(X_{1},\dots,X_{n})$.
4 Subroutines of the MCP algorithm and theoretical guarantees
In this section, we describe several key subroutines used in Algorithm 1 and prove its theoretical validity. We begin by discussing how to generate left and right p-values from the matrix of conformal p-values in Algorithm 1.